OpenStudy (anonymous):

What is the distance between point A(0, 0, 8) and B(0, 6, 0)?

OpenStudy (zzr0ck3r):

same thing

OpenStudy (raffle_snaffle):



OpenStudy (zzr0ck3r):

for P(a,b,c) and Q(d,e,f) we have \[D=\sqrt{(a-d)^2+(b-e)^2+(c-f)^2}\]

OpenStudy (zzr0ck3r):

a = 0, b = 0, c = 8, d = 0, e = 6, f = 0

OpenStudy (zzr0ck3r):

plug them in and do the arithmetic.

OpenStudy (anonymous):

2 is the distance between them

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

\[\sqrt{(0-0)^2+(0-6)^2+(8-0)^2}=\sqrt{36+64}=\sqrt{100}= ?\]

OpenStudy (zzr0ck3r):

these are vectors in 3 space, not numbers....

OpenStudy (anonymous):

10?

OpenStudy (zzr0ck3r):

correct, do you see how I did this, and now you can do the next one?

OpenStudy (anonymous):

yes i do